x square + 1 / x square -1 whole root


 

Grade:12

2 Answers

Shyam Chand
10 years ago
first of all separate as( x/x-1)+ (1/x-1)...okay...then integrate separately by taking x-1=t.. then on intgration dx=dt put it... and where there is x in numerator take x=t+1 since, x-1=t liye hue the. now solve this. beleive in yourself. that u can do it.
Shyam Chand
10 years ago
just simplify as [(x+1)/(x-1)]x[(x-1)/(x-1)] you will get simple eqn to remove under root